# Allowable Complex Black Holes in the Euclidean Gravitational Path Integral

### v1

@@ Abstract
The Euclidean Gravitational Path Integral has proven remarkably effective in the quantum regime of black hole physics. In this work, we examine the applicability of the Kontsevich-Segal-Witten (KSW) criterion for admissible complex metrics in the context of the Euclidean Gravitational Path Integral. We find that, for the super-conformal index of ${\cal N}=4$ SYM with unequal angular momenta, the black hole saddle points violate the KSW criterion precisely where the statistical description of the index breaks down. The corresponding critical point coincides with a phase transition into two-component “grey galaxy” configurations in the micro-canonical ensemble.
preprint: LITP-26-04
@@ 1 Introduction
The Euclidean Gravitational Path Integral (EGPI), formally defined in the seminal work of Gibbons and Hawking Gibbons and Hawking (1977), has historically produced results well beyond what one could expect from a semi-classical approximation. For many years, the EGPI was primarily utilized to derive the thermodynamics of black holes Hawking and Page (1983); York (1986); Whiting and York (1988); Brown and York (1993), interpreting the path integral as a partition function in the canonical ensemble where the saddle point approximation $Z\approx e^{-I_{E}}$ yields the Bekenstein-Hawking entropy. However, in recent years, this tool has experienced a resurgence Sen (2008, 2009b, 2009a); Banerjee et al. (2011); Sen (2012); Iliesiu and Turiaci (2021); Heydeman et al. (2022); Iliesiu et al. (2025a, b); Boruch et al. (2025c, e, b, d, a), proving remarkably successful in decoding the nature of quantum corrections in near-extremal and near-BPS black holes previously thought inaccessible.
A significant development occurred with the realization that the path integral must include contributions from topologies beyond the standard black hole saddle. A major leap towards resolving the Black Hole Information Paradox was achieved by considering the contribution of so-called replica wormhole saddles Almheiri et al. (2020); Penington et al. (2022), which are essential for recovering the Page curve required by unitarity. Similarly, signatures of the chaotic nature of black holes — specifically the ramp and plateau in the spectral form factor — can be recovered from the double cone geometry Saad et al. (2018); Chen et al. (2024). Yet, this success brings with it a fundamental challenge. With the understanding that novel saddle points like wormholes lead to essential new physics, the uncertainty surrounding the precise rules governing the gravitational path integral is become untenable. In particular, it is not immediately clear which saddle points should be included in the summation and which should be discarded as unphysical.
@@
Using holography, we can compute the index purely in the bulk by choosing appropriate boundary conditions determined by the chemical potentials. The corresponding bulk quantity is termed the gravitational index. The dominant saddle points contributing to the gravitational index have been identified Cabo-Bizet et al. (2019) as non-extremal analytic continuations of BPS black hole solutions. Crucially, such an analytic continuation necessarily makes the spacetime metric complex. The primary goal of this work is to test a specific proposal for the inclusion of such complex metrics: the criterion proposed by Kontsevich and Segal, and developed further by Witten (KSW) Kontsevich and Segal (2021); Witten (2021). We show that the KSW criteria are satisfied if and only if the index is well-defined.
@@
This paper is organized as follows: In Section 2, we formally define the Euclidean Gravitational Path Integral and the Superconformal Index, identifying the precise region of convergence for the SCI in the space of complex chemical potentials. In Section 2.4, we develop a practical method for implementing the KSW criterion to determine admissible complex metrics. Finally, in Section 3, we apply the KSW criterion to the complex saddle points of the gravitational index carrying unequal angular momenta and equal charges.
@@ 2.1 Euclidean Gravitational Path Integral and Saddle Points
where $I_{E}[g_{\mu\nu},\Psi]$ denotes the Euclidean action, $g_{\mu\nu}$ is the spacetime metric, and $\Psi$ represents the collection of all other dynamic fields. The partition function is contingent upon the boundary conditions specified for the fields.
@@ 2.2 The Superconformal Index
This condition ensures that states related by the action of the supercharge $\mathcal{Q}$ cancel in the trace. Consequently, the only non-vanishing contributions arise from BPS states annihilated by $\mathcal{Q}$. Therefore, the index does not depend on the variable $\beta$. In fact, due to the additional constraint (4) the SCI (3) actually depends only on four (compelx) chemical potentials.
@@
We can recover the earlier definition through the identification
@@
To evaluate the trace (5) using a path integral, we take the $CFT$ on $S^{3}\times\mathbb{R}$ and Wick rotate it to Euclidean signature by taking the time coordinate in the Lorentzian CFTCFT $t=-it_{E}$, where $t_{E}$ is the Euclidean time in the Wick rotated $CFT$. The fugacities $(\beta,\Omega_{a},\Omega_{b})$ are introduced by imposing periodic identifications on the $CFT$ coordinates,
@@
In section 3, we restrict our analysis to a special case of the superconformal index where the three chemical potentials $\Phi_{1,2,3}^{\prime}$ are taken to be equal to $\Phi^{\prime}$. The superconformal index takes the following form in this sub-sector
@@ 2.3 Convergence Conditions on the Index
This is needed to ensure that operators with large scaling dimension are suppressed in the trace, rather than enhanced. Although we are ultimately interested in the supersymmetric states for which the coefficient of $\beta$ in (3) vanishes, the cancellation of contributions from every pair of bosonic and fermionic non-supersymmetric states is properly defined only when the regulator $\beta$ satisfies the condition (13).
@@
$(E,J_{a},J_{b},Q_{1},Q_{2},Q_{3})=(1,0,0,1,0,0),\quad(1,0,0,0,1,0)\quad\text{and}\quad(1,0,0,0,0,1)\,.$ (17)
@@
The inequalities (13),(15) and (18) together bound the region of chemical potential space where the superconformal index of the $\mathcal{N}=4$ Supersymmetric Yang-Mills theory is expected to converge. The goal of this paper is to realize the boundaries of the chemical potential space through a bulk computation that identifies instabilities of the saddle points that contribute to the superconformal index.
@@ 2.4 Allowable Complex Euclidean Saddles
Direct application of the KSW conditions (19) is often computationally intractable for $p\geq 2$, but the condition can be reformulated in an equivalent form that is more practical. In a coordinate basis where the metric is diagonal $g_{ij}=\lambda_{i}\delta_{ij}$, the metric is allowable if
@@
Note that the parameters $\lambda_{i}$ are not eigenvalues of the metric in a conventional sense. Additionally, the metric transforms as a tensor $g\rightarrow J^{T}gJ$, where $J$ is the Jacobian, so the $\lambda_{i}$’s are coordinate-dependent.
To utilize (20), we must compute the diagonal entries $\lambda_{i}$ in the basis where $g_{ij}$ diagonalizes. We divide our computations into two steps. First, we enforce the $p=0,\text{ and }p=1$ conditions
@@
where $A$ and $B$ are the real and imaginary parts of the symmetric matrix $W$. The $p=1$ condition requires strictly positive real part $A>0$. Since $A$ is real and symmetric it has conventional eigenvalues, all of which must be positive.
Once the $p=1$ condition is verified, we determine whether (20) imposes further constraints. To compute the diagonal entries $\lambda_{i}$, we use the following theorem from linear algebra Horn and Johnson (1985) (Theorem 7.6.4):
@@
The diagonal entries of $K$ are the generalized eigenvalues $\kappa_{i}$ of $B$ with respect to $A$, defined by the generalized eigenvalue problem $Bx=\kappa Ax$ where the generalized eigenvector $x\neq 0$. They are determined by the characteristic equation
@@ 3 $AdS_{5}$ Black Holes with Unequal Angular Momenta
In this section we summarize the properties of the $AdS_{5}$ black hole solutions with unequal angular momenta, followed by their analytic continuations to the Euclidean signature. We focus on the non-extremal and supersymmetric configurations which dominate the gravitational index with two unequal angular fugacities. We then apply the KSW conditions to the Euclidean black hole saddles, demonstrating that they violate the criteria precisely in the region of chemical potential space where the superconformal index diverges.
@@ 3.1 Lorentzian Black Hole Solutions
The black hole solutions are presented in a frame that is static at the asymptotic boundary and with electric potential that vanishes asymptotically. These are conventional gauge choices.
@@
$q$ $=q_{*}~.$ (55)
where $r_{*}$ and $q_{*}$ refer to the abbreviations in (51-52). The real Lorentzian supersymmetric solutions therefore depend on only two parameters $(a,b)$ and can be obtained from (30) by substituting the other parameters $(m,q)$ in terms of $(a,b)$ using
@@
With this value, $r=r_{+}=r_{*}$ becomes a double root of $\Delta_{r}^{*}=0$. Therefore, the black hole is extremal. Indeed, the inverse temperature $\beta$ (43) diverges, as expected for an extremal black hole. The thermodynamic potentials (44-46) also simplify and take the values $\Omega_{a}=\Omega_{b}=\Phi=1$.
@@ 3.2 Euclidean Saddle Points for the Index
$\Delta_{r}^{*}=(r^{2}-r_{*}^{2})^{2}+\frac{1}{r^{2}}\left((r_{+}^{2}-r^{2})(1+a+b)\pm ir_{+}(r_{+}^{2}-r_{*}^{2})\right)^{2}~.$ (62)
@@
$b=\frac{r_{*}^{2}-a}{1+a}$ (63)
In the new parameterization the above conditions can be written as
$0<r_{*}<\sqrt{3},\qquad\frac{r_{*}^{2}-1}{2}<a<1,~\text{ and }~r_{+}\neq r_{*}~.$ (64)
@@
The metric for the saddle point of the index can be obtained from the Lorentzian metric for the supersymmetric black holes given in (30) in two steps. First note that the Lorentizian metric is given in a coordinate system where it becomes static at infinity i.e. asymptotically the metric becomes that of $AdS_{5}$. In such a frame, there is an ergoregion around the horizon. We define a co-rotating coordinate system : $\phi\rightarrow\phi+\Omega_{a}t$ and $\psi\rightarrow\psi+\Omega_{b}t$ so that the time coordinate $t$ is timelike everywhere outside the horizon. The analytic continuation $t\rightarrow-i\beta t_{E}$ can now be performed consistently to obtain the relevant complex Euclidean solutions whose metric now takes the form
@@
and $(m,a,b,q)$ are parameters in the solution ($\Omega_{a}$ ref.). The ranges of the coordinates are as follows: $r\in\mathbb{R}^{+},\,t_{E}\sim t_{E}+1,\,\theta\in\left[0,\frac{\pi}{2}\right],\,\phi,\psi\in\left[0,2\pi\right]$.
The determinant of the metric (70) is given by a simple expression
$\det g=\frac{\beta^{2}\rho^{4}r^{2}\sin^{2}2\theta}{4(1-a^{2})^{2}(1-b^{2})^{2}}$ (77)
In particular,
@@ 3.3 Stability of the Euclidean Saddle Points for the Index
In this section we apply the KSW criterion detailed in Section 2.4 to the complex Euclidean saddle points of the gravitational index described in Section 3.2. We show that it is both necessary and sufficient to identify the complex saddles that lie in the region where the index is well defined. The show that the instabilities first arise far from the black hole core in the plane the dominant angular momentum.
@@ 3.3.2 $p=1$ KSW condition
The simplest components of the $p=1$ conditions are the $\theta)$ and $r$ directions. First we have
@@
This is more complicated. Taking $\beta$ from (65) and $\Delta_{r}^{*}$ from (62), it is straightforward to compute ${\rm Re}\,(\beta\Delta_{r}^{*})$. The resulting expression is unwieldy, so we do not present it here. The significant feature is that it is proportional to $(r_{+}-r_{*})^{-1}$, with an overall factor that is positive at $r=r_{+}$ and a monotonically increasing function of $r$. In (84) we reported that $r_{+}-r_{*}$ and ${\rm Re}\,(\beta)$ have the same sign, and here find that ${\rm Re}\,(\beta\Delta_{r}^{*})$ has the very same sign. Therefore, the $r$ component of the $p=1$ KSW condition reduces to the $p=0$ KSW condition (82) once again.
@@
In the limit $r\rightarrow\infty$, the metric (70) can be written in the form
@@
$\frac{(1-a^{2})\sin^{2}\vartheta}{z^{2}}=(r^{2}+a^{2})\sin^{2}\theta\qquad\frac{(1-b^{2})\cos^{2}\vartheta}{z^{2}}=(r^{2}+b^{2})\cos^{2}\theta~.$ (90)
@@
This must apply for all all $\vartheta\in\left[0,\frac{\pi}{2}\right]$, which is equivalent to the two conditions,
@@
We validated our calculations numerically, with results presented in Figure 1. The green regions correspond to parameters where (95) holds at three different values of $\vartheta=0,\frac{\pi}{4},\frac{\pi}{2}$. The strongest constraints are obtained at $\vartheta=0$ and $\vartheta=\frac{\pi}{2}$. Taking the intersection of both these regions we see that the region of validity of the $p=1$ KSW condition is bounded by the curves colored red. These curves are precisely where the real parts of the renormalized potentials $\beta(1-\Omega_{a})$ or $\beta(1-\Omega_{b})$ vanish. In other words, the $p=1$ KSW criterion excludes the region where a Hilbert space interpretation of the index breaks down because the real parts of the renormalized chemical potentials $\Omega_{a,b}^{\prime}$ are negative.

### v3

@@ Abstract
The Euclidean Gravitational Path Integral has proven remarkably effective in the quantum regime of black hole physics. In this work, we examine the applicability of the Kontsevich-Segal-Witten (KSW) criterion for admissible complex metrics in the context of the Euclidean Gravitational Path Integral. We find that, for the super-conformal index of ${\cal N}=4$ SYM with unequal angular momenta, the black hole saddle points violate the KSW criterion precisely where the statistical description of the index breaks down. The corresponding critical point coincides with a phase transition into two-component “grey galaxy” configurations.
@@ 1 Introduction
The Euclidean Gravitational Path Integral (EGPI), formally defined in the seminal work of Gibbons and Hawking Gibbons and Hawking (1977), has historically produced results well beyond what one could expect from a semi-classical approximation. For many years, the EGPI was primarily utilized to derive the thermodynamics of black holes Hawking and Page (1983); York (1986); Whiting and York (1988); Brown and York (1993), interpreting the path integral as a partition function in the canonical ensemble where the saddle point approximation $Z\approx e^{-I_{E}}$ yields the Bekenstein-Hawking entropy. However, in recent years, this tool has experienced a resurgence Sen (2008, 2009b, 2009a); Banerjee et al. (2011); Sen (2012); Iliesiu and Turiaci (2021); Heydeman et al. (2022); Iliesiu et al. (2025a, b); Boruch et al. (2025c, e); Cassani et al. (2024); Boruch et al. (2025b, d); Cassani et al. (2025); Boruch et al. (2025a), proving remarkably successful in decoding the nature of quantum corrections in near-extremal and near-BPS black holes previously thought inaccessible.
A significant development occurred with the realization that the path integral must include contributions from topologies beyond the standard black hole saddle. Major progress towards resolving the Black Hole Information Paradox was achieved by considering the contribution of so-called replica wormhole saddles Almheiri et al. (2020); Penington et al. (2022), which are essential for recovering the Page curve required by unitarity. Similarly, signatures of the chaotic nature of black holes — specifically the ramp and plateau in the spectral form factor — can be recovered from the double cone geometry Saad et al. (2018); Chen et al. (2024). Yet, this success brings with it a fundamental challenge. With the understanding that novel saddle points like wormholes lead to essential new physics, the uncertainty surrounding the precise rules governing the gravitational path integral becomes untenable. In particular, it is not immediately clear which saddle points should be included in the summation, and which should be discarded as unphysical.
@@
Using holography, we can compute the index purely in the bulk by choosing appropriate boundary conditions determined by the chemical potentials. The corresponding bulk quantity is termed the gravitational index. The dominant saddle points contributing to the gravitational index have been identified as non-extremal analytic continuations of BPS black hole solutions Cabo-Bizet et al. (2019) . Crucially, such an analytic continuation necessarily makes the spacetime metric complex. The primary goal of this work is to test a specific proposal for the inclusion of such complex metrics: the criterion proposed by Kontsevich and Segal, and developed further by Witten (KSW) Kontsevich and Segal (2021); Witten (2021). We show that the KSW criteria are satisfied precisely if and only if the index is well-defined.
@@
This paper is organized as follows: In Section 2, we introduce the Euclidean Gravitational Path Integral and the Superconformal Index, identifying the precise region of convergence for the SCI in the space of complex chemical potentials. In Section 2.4 we develop a practical method for implementing the KSW criterion to determine admissible complex metrics. In Section 3 we introduce the Lorentzian and Euclidean versions of the AdS5 black holes with two independent angular momenta and a single electric charge. We identify the solutions that preserve supersymmetry and the subset of those that satisfy the KSW conditions. Those are the allowed complex saddle points that contribute of the gravitational index. Finally, in Section 4, we compare to other recent work and discuss open problems.
@@ 2.1 Euclidean Gravitational Path Integral and Saddle Points
where $I_{E}[g_{\mu\nu},\Psi]$ denotes the Euclidean action, $g_{\mu\nu}$ is the spacetime metric, and $\Psi$ represents the collection of all other dynamic fields. All fields are subject to boundary conditions.
@@ 2.2 The Superconformal Index
This condition ensures that states related by the action of the supercharge $\mathcal{Q}$ cancel in the trace. Consequently, all non-vanishing contributions arise from BPS states annihilated by $\mathcal{Q}$. Therefore, the index does not depend on the variable $\beta$. In fact, due to the additional constraint (4) the SCI (3) actually depends only on four (complex) chemical potentials.
@@
We can recover the earlier definition (3) through the identifications
@@
To evaluate the trace (5) using a path integral, we consider the $CFT$ on $S^{3}\times\mathbb{R}$ and Wick rotate it to Euclidean signature by taking the time coordinate in the Lorentzian CFTCFT $t=-it_{E}$, where $t_{E}$ is the Euclidean time in the Wick rotated $CFT$. The fugacities $(\beta,\Omega_{a},\Omega_{b})$ are introduced by imposing periodic identifications on the $CFT$ coordinates,
@@
In section 3, we restrict our analysis to the special case of the superconformal index where the three chemical potentials $\Phi_{1,2,3}^{\prime}$ are taken to be equal and denoted $\Phi^{\prime}$. In this sub-sector, the superconformal index is
@@ 2.3 Convergence Conditions on the Index
This is needed to ensure that operators with large scaling dimension are suppressed in the trace, rather than enhanced. Although we are ultimately interested in supersymmetric states, for which the coefficient of $\beta$ in (3) vanishes, the cancellation of contributions from every pair of bosonic and fermionic non-supersymmetric states is properly defined only when the regulator $\beta$ satisfies the condition (13).
@@
$(E,J_{a},J_{b},Q_{1},Q_{2},Q_{3})=(1,0,0,1,0,0)~,\quad(1,0,0,0,1,0)~<\quad\text{and}\quad(1,0,0,0,0,1)\,.$ (17)
@@
The inequalities (13), (15) and (18) together bound the region of chemical potential space where the superconformal index of the $\mathcal{N}=4$ Supersymmetric Yang-Mills theory is expected to converge. The goal of this paper is to realize these boundaries through a bulk computation that identifies instabilities of the saddle points that contribute to the superconformal index.
@@ 2.4 Allowable Complex Euclidean Saddles
Direct application of the KSW conditions (19) is often impractical for $p\geq 2$, but the condition can be reformulated in an equivalent form that is more convenient. In a coordinate basis where the metric is diagonal $g_{ij}=\lambda_{i}\delta_{ij}$, the metric is allowable if
@@
Note that the parameters $\lambda_{i}$ are not eigenvalues of the metric in a conventional sense. The metric is symmetric but, because it is complex, it is not Hermitean. Additionally, the metric transforms as a tensor $g\rightarrow J^{T}gJ$, where $J$ is the Jacobian, so the $\lambda_{i}$’s are coordinate-dependent.
To utilize (20), we must compute the diagonal entries $\lambda_{i}$ in the basis where $g_{ij}$ diagonalizes [^2] See also Appendix B of Chryssanthacopoulos .. We divide our computations into two steps. First, we enforce the $p=0,\text{ and }p=1$ conditions
@@
where $A$ and $B$ are the real and imaginary parts of the symmetric matrix $W$. The $p=1$ condition requires strictly positive real part $A>0$. Since $A$ is real and symmetric it has conventional eigenvalues which, according to the KSW condition, must be positive.
Once the $p=0$ and $p=1$ conditions are verified, we determine whether (20) imposes further constraints. To compute the diagonal entries $\lambda_{i}$, we use the following theorem from linear algebra (Theorem 7.6.4 in Horn and Johnson (1985)):
@@
The diagonal entries of $K$ are the generalized eigenvalues $\kappa_{i}$ of $B$ with respect to $A$, defined as solutions to the generalized eigenvalue problem $Bx=\kappa Ax$ where the generalized eigenvector $x\neq 0$. They are determined by the characteristic equation
@@ 3 $AdS_{5}$ Black Holes with Unequal Angular Momenta
In this section we summarize the properties of the $AdS_{5}$ black hole solutions with unequal angular momenta, followed by their analytic continuations to the Euclidean signature. We focus on the non-extremal, supersymmetric configurations which dominate the gravitational index. We then apply the KSW conditions to the Euclidean black hole saddles, and show that they violate the criteria precisely in the region of chemical potential space where the superconformal index diverges.
@@ 3.1 Lorentzian Black Hole Solutions
The black hole solutions are presented in a frame that is static at the asymptotic boundary and with electric potential that vanishes asymptotically. These are the conventional gauge choices.
@@
$q$ $=q_{*}~,$ (55)
where $r_{*}$ and $q_{*}$ refer to the abbreviations in (51-52). The real Lorentzian supersymmetric solutions therefore depend on only two parameters $(a,b)$ and can be obtained from (30) by eliminating the other parameters $(m,q)$ in terms of $(a,b)$ using
@@
With this value, $r=r_{+}=r_{*}$ becomes a double root of $\Delta_{r}^{*}=0$. Therefore, the black hole is extremal. Indeed, the inverse temperature $\beta$ (43) diverges, as expected for an extremal black hole. The thermodynamic potentials (44-46) also simplify and take the values $\Omega^{*}_{a}=\Omega^{*}_{b}=\Phi^{*}=1$.
@@ 3.2 Euclidean Saddle Points for the Index
$\Delta_{r}^{*}=(r^{2}-r_{*}^{2})^{2}+\frac{1}{r^{2}}\Big((r_{+}^{2}-r^{2})(1+a+b)\pm ir_{+}(r_{+}^{2}-r_{*}^{2})\Big)^{2}~.$ (62)
@@
$b=\frac{r_{*}^{2}-a}{1+a}~.$ (63)
In the new parameterization the inequalities (61) can be written as
$0<r_{*}<\sqrt{3}~,\qquad\frac{r_{*}^{2}-1}{2}<a<1~.$ (64)
@@
The metric for the saddle point of the index can be obtained from the Lorentzian metric for the supersymmetric black holes given in (30) in two steps. First recall that the Lorentizian metric is given in a coordinate system where it becomes static at infinity i.e. asymptotically the metric becomes that of $AdS_{5}$. In such a frame, there is an ergoregion around the horizon. The replacements $\phi\rightarrow\phi+\Omega_{a}t$ and $\psi\rightarrow\psi+\Omega_{b}t$ with $t$ fixed define a co-rotating coordinate system so that the time coordinate $t$ is timelike everywhere outside the horizon. Then the analytic continuation $t\rightarrow-i\beta t_{E}$ can be performed to obtain the complex Euclidean solutions with metric
@@
and $(m,a,b,q)$ are parameters in the solution and the quantities $(\beta,\Omega_{a},\Omega_{b})$ are defined in (65)-(67). The ranges of the coordinates are: $r\in\mathbb{R}^{+},\,t_{E}\sim t_{E}+1,\,\theta\in\left[0,\frac{\pi}{2}\right],\,\phi,\psi\in\left[0,2\pi\right]$.
The metric (70) has simple subdeterminants:
@@
The full determinant becomes:
$\det g=\frac{\beta^{2}\rho^{4}r^{2}\sin^{2}2\theta}{4(1-a^{2})^{2}(1-b^{2})^{2}}~.$ (79)
@@ 3.3 Stability of the Euclidean Saddle Points for the Index
In this section we apply the KSW criterion detailed in Section 2.4 to the complex Euclidean saddle points of the gravitational index described in Section 3.2. We show that it is both necessary and sufficient to identify the complex saddles that lie in the region where the index is well defined. The show that the instabilities first arise far from the black hole core, in the plane the dominant angular momentum.
@@ 3.3.2 $p=1$ KSW condition
The simplest components of the $p=1$ conditions are the $\theta$ and $r$ directions. First we have
@@
This is more complicated. Taking $\beta$ from (65) and $\Delta_{r}^{*}$ from (62), it is straightforward to compute ${\rm Re}\,(\beta\Delta_{r}^{*})$. The resulting expression is unwieldy, so we do not present it here. The significant feature is that it is proportional to $(r_{+}-r_{*})^{-1}$, with an overall factor that is positive at $r=r_{+}$ and a monotonically increasing function of $r$. After (83) we concluded that $r_{+}-r_{*}$ and ${\rm Re}\,(\beta)$ have the same sign, and here find that ${\rm Re}\,(\beta\Delta_{r}^{*})$ has the very same sign. Therefore, the $r$ component of the $p=1$ KSW condition reduces to the $p=0$ KSW condition (82) once again.
@@
In the asymptotic region $r\rightarrow\infty$, the metric (70) can be written in the form
@@
$\frac{(1-a^{2})\sin^{2}\vartheta}{z^{2}}=(r^{2}+a^{2})\sin^{2}\theta~,\qquad\frac{(1-b^{2})\cos^{2}\vartheta}{z^{2}}=(r^{2}+b^{2})\cos^{2}\theta~.$ (90)
@@
This equation must apply for all all $\vartheta\in\left[0,\frac{\pi}{2}\right]$, so it is equivalent to the two conditions,
@@
We validated our calculations numerically, with results presented in Figure 1. The three real parameters $(a,r_{+},r_{*})$ on configuration space are visualized in plots as the $(a,r_{+})$ cross-section at a fixed value of $r_{*}$. At fixed $r_{*}$, the range of the parameter $r_{+}$ is $r_{+}>r_{*}$, because of the $p=0$ condition (84), and the parameter $a$ takes values between $a_{\text{min}}=(r_{*}^{2}-1)/2$ and $a_{\text{max}}=1$, as shown in (64). We picked the value $r_{*}=0.1$ in all plots. The green regions in the three plots correspond to parameters where (95) holds at three different values of $\vartheta=0,\frac{\pi}{4},\frac{\pi}{2}$. The strongest constraints are obtained at $\vartheta=0$ and $\vartheta=\frac{\pi}{2}$. Taking the intersection of both these regions we see that the region of validity of the $p=1$ KSW condition is bounded by the curves colored red. These curves are precisely where the real parts of the renormalized potentials $\beta(1-\Omega_{a})$ or $\beta(1-\Omega_{b})$ vanish. In other words, the $p=1$ KSW criterion excludes the region where a Hilbert space interpretation of the index breaks down because the real parts of the renormalized chemical potentials $\Omega_{a,b}^{\prime}$ are negative.